OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1 + 4 + 7 + . . . + (3n - 2) = n(3n - 1)/2

OpenStudy (anonymous):

2. An artist is creating a mosaic that cannot be larger than the space allotted which is 4 feet tall and 6 feet wide. The mosaic must be at least 3 feet tall and 5 feet wide. The tiles in the mosaic have words written on them and the artist wants the words to all be horizontal in the final mosaic. The word tiles come in two sizes: The smaller tiles are 4 inches tall and 4 inches wide, while the large tiles are 6 inches tall and 12 inches wide. If the small tiles cost $3.50 each and the larger tiles cost $4.50 each, how many of each should be used to minimize the cost? What is the minimum cost? Show work

OpenStudy (aravindg):

Post a single question at a time.

OpenStudy (anonymous):

sorry about that

ganeshie8 (ganeshie8):

for Sk, simply replace n wid k in the given statement : Sn: 1 + 4 + 7 + . . . + (3n - 2) = n(3n - 1)/2 Sk: 1 + 4 + 7 + . . . + (3k - 2) = k(3k - 1)/2

ganeshie8 (ganeshie8):

for Sk+1, replace n wid (k+1) and simplify if possible

OpenStudy (anonymous):

my teacher told me to replace it with n though

OpenStudy (anonymous):

(k+1) & sk+1

ganeshie8 (ganeshie8):

-.-